OpenStudy (anonymous):

You have 57.0ml of a 0.400M stock solution that must be diluted to 0.100M. Assume volumes are addictive. How much water should you add?

OpenStudy (aaronq):

use \(M_1V_1=M_2V_2\) then add water necessary to achieved the req'd volume.

OpenStudy (anonymous):

What?

OpenStudy (aaronq):

use the dilution formula (above), once you find what the volume should be, adjust it by adding the amount of water necessary.

OpenStudy (anonymous):

does that give you 228?

OpenStudy (aaronq):

units?

OpenStudy (anonymous):

288M

OpenStudy (anonymous):

228m

OpenStudy (aaronq):

they're asking for the amount of water you should add. it should be in units of volume.

OpenStudy (anonymous):

so 0.288

OpenStudy (aaronq):

it's the correct value, given you use the appropriate units.

OpenStudy (anonymous):

Right

OpenStudy (aaronq):

but note that that is the TOTAL volume and you already have 57 mL

OpenStudy (anonymous):

Ok so... it's 0.228M

OpenStudy (aaronq):

M is in units of concentration.

OpenStudy (aaronq):

M = mol/L

OpenStudy (anonymous):

Ok i'm getting confused.

OpenStudy (aaronq):

you need units of volume. \(M_1V_1=M_2V_2 \rightarrow 0.057L*0.4 M=V_2*0.1 M\) \(V_2=\dfrac{0.057L*0.4 M}{0.1 M}=0.288 L\) but you already have 57 mL. 0.288 L - 0.057 L =your answer

OpenStudy (anonymous):

Oh... I have to subtract it from 0.0057

OpenStudy (aaronq):

you subtract 0.057 from it, yes. because you're adding water to it to reach that volume

OpenStudy (anonymous):

Makes sense now. Thank you.